OpenStudy (anonymous):

If G(x) = (x + 4)2, and F(x) = x2 + 8x, what is the function G(F(x))?

OpenStudy (anonymous):

is it G(x) = (x+4)^2, or (x+4)*2?

OpenStudy (anonymous):

i'd say you meant (x+4)^2==>G(F(x))=g(x^2+8x)=(x^2+8x+4)^2

OpenStudy (anonymous):

x^2